OpenStudy (anonymous):

find exact value of tan -1(1)

OpenStudy (xapproachesinfinity):

well think about the inverse of arctan which is tan. now that angle measure make tan 1

OpenStudy (xapproachesinfinity):

tan(?)=1

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

for that we look at this special triangle that has two side 1 and two angles 45 we know that tan=adjacent/hypotenuse tan(45)=1/1=1 so the angle is 45 degrees now going back the what we said arctan is inverse of tan so if we have tan(45)=1===> 45=arctan(1) mean that arctan(1)=45 (pi/4 is radians)